According to the NEC Table 220.55, what would the demand load be for two 14 kW ranges?
Answer #1

According to NEC Table 220.55 ==>

Demand load for 2, 12 kW range = D2,12 = 11 kW (See column - C)

Therefore, Demand load for 2, 14 kW range = (1+0.05*(14-12))*D2,12 = (1+0.10)*11 = 1.1*11 = 12.1 kW ==> ANSWER